OpenStudy (stephy321):

Hi. So, I've been having difficulty with the chain rule for the past week. I haven't asked my teacher for help because his way of teaching is extremely confusing and not helpful at all. I would appreciate if someone could walk me through the problem below. Thank you.

OpenStudy (sshayer):

post your problem

OpenStudy (stephy321):

here is the problem

OpenStudy (sshayer):

i can't read properly

OpenStudy (sshayer):

is it \[\frac{ 1 }{ 2 }x^2\sqrt{16-x^2}\] ?

OpenStudy (sshayer):

is it correct?

OpenStudy (sshayer):

are you here?

OpenStudy (stephy321):

it equals y. but other than that, it is correct

OpenStudy (sshayer):

you want to find dy/dx.

OpenStudy (sshayer):

i will be back after dinner.wait for few minutes.

OpenStudy (stephy321):

okay, I will be here

OpenStudy (sshayer):

\[\frac{ dy }{ uv }=uv \prime+u \prime v\] where u and v are functions of x

OpenStudy (sshayer):

\[\frac{ dy }{ dx }=\frac{ d }{ dx }\left[ \frac{ 1 }{ 2 }x^2\left( 16-x^2 \right)^{\frac{ 1 }{ 2 }} \right]\] \[=\frac{ 1 }{ 2 }\left[ x^2 \times \frac{ 1 }{ 2 }\left( 16-x^2 \right)^{\frac{ 1 }{ 2 }-1}\left(0 -2x \right)+2x \left( 16-x^2 \right)^{\frac{ 1 }{ 2 }} \right]\] \[=\frac{ 1 }{ 2 }\left[ \frac{ -x^3 }{ \sqrt{16-x^2} }+2x \sqrt{16-x^2} \right]\] \[=\frac{ 1 }{ 2 }\left[ \frac{-x^3+2x \left( 16-x^2 \right) }{ \sqrt{16-x^2} } \right]\] \[=\frac{ 1 }{ 2 }\left[ \frac{ -x^3+32x-2x^3 }{ \sqrt{16-x^2} } \right]\] \[=\frac{ 32x-3 x^3 }{2 \sqrt{16-x^2} }\]

OpenStudy (sshayer):

i am leaving now

OpenStudy (stephy321):

Thank you for your help

OpenStudy (sshayer):

you have any questions.